Problem#1(40 Points) An urn contains two balls, one is red and one is black. A ball is drawn at random and its color is noted then returned back to the urn. This process is repeated, and stoped when the red ball is drawn or until a maximum of 6 trials. Let the random variable X denotes the number of trials. 1. Find the p.m.fof X. (10 Points)
Answer #1

1) below is pmf of X:

P(X=1) =P(first ball is red) =1/2

P(X=2) =P(first ball not red and second ball is red) =(1/2)*(1/2) =1/4

P(X=3) =P(first two ball not red and second ball is red) =(1/2)2*(1/2) =1/8

P(X=4 )=P(first three ball not red and second ball is red) =(1/2)3*(1/2) =1/16

P(X=5) =P(first four ball not red and second ball is red) =(1/2)4*(1/2) =1/32

P(X=6) =P(first 5 ball are not red) =(1/2)5 =1/32

hence below is pmf of X:

x f(x)
1    1/2
2    1/4
3    1/8
4    1/16
5    1/32
6    1/32